OpenStudy (anonymous):

What is (f ◦ g)(x) of f(x) = 5x - 26 and g(x) = 6x

jimthompson5910 (jim_thompson5910):

(f ◦ g)(x) = f(g(x)) f(x) = 5x - 26 f(g(x)) = 5(g(x)) - 26 f(g(x)) = 5(6x) - 26 f(g(x)) = 30x - 26 So (f ◦ g)(x) = 30x - 26